OpenStudy (cavaliers999):

A student is running at her top speed of 5.2 m/s to catch a bus, which is stopped at the bus stop. When the student is still a distance 41.5 m from the bus, it starts to pull away, moving with a constant acceleration of 0.171 m/s2 . What is the minimum speed the student must have to just catch up with the bus?

OpenStudy (solomonzelman):

((The way I will do it, is through calculus.)) Function for the bus displacement: (41.5 is the initial distance of bus from student) \(\color{#0000ff }{ \displaystyle {\rm x}_1(t)=(0.171/2)t^2+41.5}\) Student's function for displacement: (We don't know the velocity the student goes at when bus starts to pull away from the bus-stop) \(\color{#0000ff }{ \displaystyle {\rm x}_2(t)={\bf v}{\tiny~}t}\) When can the functions intersect? \(\color{#0000ff }{ \displaystyle {\rm x}_1(t)={\rm x}_2(t)}\) \(\color{#0000ff }{ \displaystyle 0.0855{\tiny~}t^2+41.5={\bf v}{\tiny~}t}\) \(\color{#0000ff }{ \displaystyle {\bf v}=0.0855t+(41.5/t)}\) Let's minimize this function! Recall the constraints given in the problem, and by definition of "time". \(\color{#0000ff }{ \displaystyle 0\le {\bf v}\le 5.2~~{m/s}}\) and \(\color{#0000ff }{ \displaystyle t>0~~{s}}\) I will minimize the function: \(\color{#0000ff }{ \displaystyle {\bf v}'=0.0855-(41.5/t^2)}\) the only critical number is: \(\color{#0000ff }{ \displaystyle t=22~m/s}\) (this is the time when min velocity is reached) \(\color{#0000ff }{ \displaystyle {\bf v}_{\rm min}={\bf v}(22)=3.77~m/s}\) So, I get that \(\color{#ff0000}{ \displaystyle {\bf v}_{\rm min}=3.77~m/s}\) as the answer.

OpenStudy (cavaliers999):

how did you get the function for the bus displacement? I'm a bit confused about that thanks